OpenStudy (anonymous):

Help!!! Integrate ln | 1+x |

terenzreignz (terenzreignz):

\[\Large \int \ln|1+x|dx\] this?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Well then, all you need for this is to recall the way to integrate *by parts* and the derivative of the natural logarithm. INTEGRATION BY PARTS:\[\Large \int udv = uv - \int vdu\] \[\Large \frac{d}{dt}\ln|t| = \frac1t\]

OpenStudy (anonymous):

how to integrate by parts if there is only one part?

terenzreignz (terenzreignz):

Good question... much like how the skin is so... second nature, people don't usually call it an organ like, say, the liver, the stomach, etc But I digress ^_^ THESE are two parts in themselves... people hardly seem to count the differential dx as a part in itself :P \[\Large \int \color{green}{\ln|1+x|}\color{red}{dx}\]

terenzreignz (terenzreignz):

So let's begin. Let u = ln|1+x| and dv = dx Can you find du and v?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

du= 1/1+x v=x

OpenStudy (anonymous):

is it?

terenzreignz (terenzreignz):

That is correct :) Apply the formula: \[\Large = x\ln|1+x| - \color{blue}{\int \frac{x}{1+x}dx}\] I trust that 'blue' integral won't be a problem? ^_^

OpenStudy (anonymous):

for da blue part do integration by parts too??

terenzreignz (terenzreignz):

Nope. \[\Large \frac{x}{1+x}= \frac{x\color{red}{+1-1}}{1+x}= \frac{x+1}{x+1}-\frac1{1+x}=1-\frac1{1+x}\] \[\Large \implies \int\frac{x}{1+x}dx=\int\left[1-\frac1{1+x}\right]dx\] Much easier to integrate, no? ^_^

terenzreignz (terenzreignz):

Much of the time, integration calls for creativity, so keep this technique in mind, and use it if the situation calls for it ^_^

OpenStudy (anonymous):

how can we just put +1 n -1 like tat?

terenzreignz (terenzreignz):

We can... we basically added zero ^_^ +1-1 = 0 So it does nothing, except give us an *equivalent* expression which just happened to be easier to integrate. Such tricks come a plenty in maths, so watch out for them. Look, I need to go now, I'm sure someone can pick up where we left off... Good luck ^_^ ----------------------------------- Terence out :)

OpenStudy (anonymous):

:'( Thanks btw i'll try my best